OpenStudy (anonymous):

Find the limit of the (An)n>=1 A(n)=f(1)*f(2)*...*f(n) f:R\{-3/2}->R;f(x)=(x+1)/(2x+3)

OpenStudy (anonymous):

I've used the Stoltz-Cesaro theorem and gave me the value 1/2 But my answer is wrong

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Since \[ \lim_{n->\infty} f(n)=\frac 1 2 \] There is exist N so n>N \[ 0< f(n) < \frac 3 4\\ 0\le A(n)\le f(1)*f(2)* \cdots f(N)* f(N+1) *fN(+2)*\cdots f(n)\le \\ f(1)*f(2)* \cdots f(N)\left ( \frac 34 \right)^{n-N}=C_N \left ( \frac 34 \right)^{n-N} \] If n goes to infinity, then A(n) goes to zero, since \[ \lim_{n->\infty} C_N \left ( \frac 34 \right)^{n-N}=0 \]